OpenStudy (anonymous):

Algebra 1 Help Please! Create your own equation written in standard form. Determine any point that is a solution and will be on this line. Justify your answer algebraically. Explain how to find the x- and y- intercept(s) using complete sentences. Write the intercepts as ordered pairs. Rewrite the equation using function notation. Graph the function and show the intercepts.

OpenStudy (anonymous):

HI :DDD

OpenStudy (anonymous):

lol That's not the answer! :p

OpenStudy (anonymous):

HI again :DDD

OpenStudy (anonymous):

Well are you able to come up with an equation?

OpenStudy (anonymous):

2x+7y=14 ?

OpenStudy (agent0smith):

Explain how to find the x- and y- intercept(s) using complete sentences. ^ to find x-int, plug in y=0 to find y-int, plug in x=0

OpenStudy (anonymous):

x=1 y=1?

OpenStudy (agent0smith):

no... 2x+7y=14 ^ to find x-int, plug in y=0. Then, to find y-int, plug in x=0.

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

Sorry about that!

OpenStudy (agent0smith):

But you forgot "Write the intercepts as ordered pairs." Do them one at a time.

OpenStudy (anonymous):

(7,2)

OpenStudy (agent0smith):

I was going to warn you to not make that mistake. ^ to find x-int, plug in y=0... your x-intercept is x=2. Write it as an ordered pair.

OpenStudy (anonymous):

(2,0) (0,7) ?

OpenStudy (agent0smith):

Much better.

OpenStudy (agent0smith):

Notice that an x intercept can NEVER EVER EVER be (7,2), and neither can a y-intercept. Why?

OpenStudy (anonymous):

Thank you! Because it can only be an x intercept and if you put (7,2) its a y intercept and an x intercept?

OpenStudy (agent0smith):

An x-intercept has to be on the x-axis, just like a y-intercept has to be on the y-axis. (7,2) is never any kind of intercept. Get a graph, plot a point at (7,2) and see why.

OpenStudy (agent0smith):

Or, for a point to be an x-intercept, what must y be equal to?

OpenStudy (anonymous):

The point isnt on the x-axis? and the y must be 0?

OpenStudy (agent0smith):

Exactly. What about for a y-intercept... What does x have to?

OpenStudy (anonymous):

the x has to be a zero!

OpenStudy (agent0smith):

Exactly. Learn this now, because I have algebra 2 students who don't understand this, and it makes me want to murder them :P

OpenStudy (anonymous):

haha! Okay, I will try to remember that!

OpenStudy (agent0smith):

You better! Because if you somehow end up in my algebra 2 class, and complain to the cops that your algebra 2 teacher murdered you, I will point them to this and they'll say "Well, he did warn you... ". Cops probably don't listen to dead people anyway :P

OpenStudy (anonymous):

hahaha! wow okay i will write it on a sticky note and post it to my forehead!!

OpenStudy (agent0smith):

Hahaha okay :) And don't worry I wouldn't actually murder you :P but my god it is frustrating to have algebra 2 students, who don't even understand the basics of algebra 1.

OpenStudy (anonymous):

lol I understand completely! I will make sure to go into Algebra 2, fully informed of this problem! Thank you soooo much for helping me on this problem! I have gymnastics practice so I have to go! I appreciate you helping me!

OpenStudy (agent0smith):

You're welcome! Good job :)